OpenStudy (anonymous):

The sum of w and 54 is equal to s. Let s = 72. Which equation can be used to find the value of w?

OpenStudy (anonymous):

Well, the first sentence suggests that w and s are related as follows: \[w + 54 = s\]

OpenStudy (anonymous):

So, if s = 72, then your equation becomes w + 54 = 72. Solve for w.

OpenStudy (anonymous):

W+54=s ....eq1 S=72 ......eq2 by substitution of eq2 in eq1 we get w+54=72 ==> w=72-54 ==>w= 18